OpenStudy (anonymous):

What are the roots of the equation x2 + 5x − 3 = 0?

OpenStudy (anonymous):

x = 1/5

OpenStudy (solomonzelman):

\[x^2 + 5x − 3 = 0\] Quadratic formula would be the best,

OpenStudy (e.mccormick):

@freakenpurplepizza not sure where you got that, but that is not it.

OpenStudy (e.mccormick):

@tori67 SolomonZelman is pointing you the right direction.

OpenStudy (solomonzelman):

\[\huge\color{blue}{x~~=~~~ \frac{-b±\sqrt{b^2-4ac}}{2a} } \]

OpenStudy (solomonzelman):

\[\huge\color{blue}{1} x^2+\huge\color{blue}{5} x-\huge\color{blue}{3} =0\] compare to,\[\huge\color{blue}{a} x^2+\huge\color{blue}{b} x+\huge\color{blue}{c} =0\]